OpenStudy (anonymous):

What are the solutions of 5x2 = -15x?

OpenStudy (anonymous):

5x^2 = -15x bring everything to one side 5x^2 + 15x = 0 factor 5x(x+3) = 0 so what are the roots of the equation?

OpenStudy (anonymous):

aka what values for x satisfy the equation?

OpenStudy (anonymous):

5x2 + 15x

OpenStudy (anonymous):

You could do it that way or use the quadratic equation.

OpenStudy (anonymous):

this is a challenge question on my math packet and i have no idea how to do it

OpenStudy (anonymous):